OpenStudy (anonymous):

write a product of the form (a+√b)(a-√b) so that the answer, when simplified, is 59.

myininaya (myininaya):

(8+sqrt(5))(8-sqrt(5))=64-5=59

OpenStudy (anonymous):

thanks

myininaya (myininaya):

do you want to know how i got that

OpenStudy (anonymous):

sure its so kind of u

myininaya (myininaya):

(a+sqrt(b))(a-sqrt(b))=a^2-b correct?

myininaya (myininaya):

we want a^2-b=59

myininaya (myininaya):

so I added b on both sides

myininaya (myininaya):

so we have a^2=59+b and I just chose b to make the right hand side a perfect sqare

myininaya (myininaya):

I let b=5 a^2=59+5 a^2=64 a=8

myininaya (myininaya):

you didn't have to chose b to be 5 though

myininaya (myininaya):

you could let b be anything and then find a the same away I did above

myininaya (myininaya):

any questions?